# Informed Asymmetric Actor-Critic: Leveraging Privileged Signals Beyond Full-State Access

### v1

@@ Abstract
Reinforcement learning in partially observable environments requires agents to act under uncertainty from noisy, incomplete observations. Asymmetric actor-critic methods leverage privileged information during training to improve learning under these conditions. However, existing approaches typically assume full-state access during training. In this work, we challenge this assumption by proposing a novel actor-critic framework, called informed asymmetric actor-critic, that enables conditioning the critic on arbitrary privileged signals without requiring access to the full state. We show that policy gradients remain unbiased under this formulation, extending the theoretical foundation of asymmetric methods to the more general case of privileged partial information. To quantify the impact of such signals, we propose informativeness measures based on kernel methods and return prediction error, providing practical tools for evaluating training-time signals. We validate our approach empirically on benchmark navigation tasks and synthetic partially observable environments, showing that our informed asymmetric method improves learning efficiency and value estimation when informative privileged inputs are available. Our findings challenge the necessity of full-state access and open new directions for designing asymmetric reinforcement learning methods that are both practical and theoretically sound.
@@ 1 INTRODUCTION
Reinforcement learning (RL) has emerged as a powerful tool for optimizing control policies in various domains, including the control of heating, ventilation, and air conditioning systems [1], energy systems [2, 3], autonomous driving [4], and robotics [5]. However, when deploying RL in such real-world applications, agents must often operate under partial observability, relying on incomplete and noisy observations to make decisions. This setting is formalized by partially observable Markov decision processes (POMDPs) [6], where optimal actions depend on the history of past observations and actions.
@@
To address this, RL methods for fully observable settings have been adapted by learning history-dependent policies, typically using recurrent neural networks (RNNs) to encode observation-action sequences [7, 8]. Although these methods are, in principle, capable of learning optimal history-dependent policies, they assume the same level of observability during training and execution, constraining policy learning to the limited information available at deployment. Yet, in practice, this assumption is unnecessarily restrictive and possibly suboptimal. Many training environments provide privileged information unavailable at execution, such as diagnostic sensors or simulators exposing internal variables, without necessarily providing full access to the true state. Leveraging such asymmetric observability motivates the paradigm of asymmetric learning, which aims to exploit additional information at training while ensuring deployable history-dependent policies.
Asymmetric actor-critic methods provide a framework for this setting, as they condition the actor on observable histories while allowing the critic access to privileged information during training [9, 10]. However, existing approaches often assume either full-state access or no additional information, leaving the more general case of privileged partial information underexplored.
In this work, we introduce the informed asymmetric actor-critic framework, which generalizes prior asymmetric approaches by allowing the critic to condition on arbitrary state-dependent privileged inputs, without requiring full-state access. We show that this formulation yields unbiased estimators of value functions and policy gradients, extending the theoretical foundation of asymmetric learning to a broader class of training-time signals. To guide the selection and evaluation of privileged inputs, we propose two informativeness criteria: (i) a pre-training measure based on the Hilbert-Schmidt Conditional Independence Criterion (HSCIC), and (ii) a post-training metric based on return prediction error. These tools allow practitioners to quantify the utility of privileged signals before or after policy optimization. We empirically validate our informed asymmetric method on benchmark navigation tasks and in synthetic informed POMDP environments, demonstrating improved policy learning when informative privileged signals are used. Our findings highlight the importance of informativeness and challenge the assumption that full-state access is essential for asymmetric reinforcement learning.
2 RELATED WORK
Traditional RL methods have been adapted to partially observable settings by learning history-dependent policies that process sequences of past observations and actions using RNNs [11, 8, 7, 12, 13]. Since directly optimizing from raw histories is challenging, many approaches compress these sequences into compact latent representations, often by introducing auxiliary learning objectives [14, 15, 16].
@@ 2 RELATED WORK
Some works address partial observability by training privileged expert policies conditioned on true states and imitating them [17]. However, these methods often lack theoretical guarantees and may lead to suboptimal policies in POMDPs [18]. To mitigate this, Warrington et al. [18] propose constraining the expert policy to yield an optimal policy under partial observability. Another line of work exploits privileged information in model-based RL by constructing world models that summarize histories or integrate additional state signals. Examples include the Informed Dreamer [19], the Wasserstein Believer [20], and the Scaffolder [21].
Asymmetric actor-critic methods have emerged as a simple yet powerful framework for leveraging privileged information during training. By conditioning the critic on the full state and the actor on the history, these methods aim to guide policy updates more effectively. Pinto et al. [9] introduce an early asymmetric actor-critic approach that achieves strong empirical performance but suffers from biased gradient estimates [10]. Baisero and Amato [10] address this issue by introducing the history-state value function, explicitly modeling the relationship between histories and latent states to ensure unbiased gradients.
Recent theoretical work has established convergence guarantees for policy gradient and actor-critic methods in both fully and partially observable settings, including symmetric recurrent natural actor-critic methods using RNNs [14, 22] and asymmetric settings with linear function approximators [23].
Despite these advances, existing actor-critic methods typically assume either full access to the true state during training or no additional information at all. However, many real-world settings fall between these extremes: some internal variables may be observable during training, while others remain hidden or only partially measurable. Methods that exploit privileged partial information, without requiring full-state access, are vastly unstudied, and it will be the main focus of this article.
@@ 3 BACKGROUND
In this section, we introduce the formal notion of partially observable Markov decision processes (POMDPs) and the informed POMDP framework, which motivates our informed asymmetric actor-critic framework.
@@ 3.1 Partially Observable Markov Decision Processes
A partially observable Markov decision process (POMDP) [6] models sequential decision-making under uncertainty as tuple $(\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,P,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{O}$ denote the state, action, and observation spaces. The transition probabilities $T(s^{\prime}\mid s,a)$ describe the process dynamics. The agent emits observations via $O(o\mid s)$ and selects actions $a_{t}$ based on the observable history $h$, defined as the sequence of past observations and actions. It receives a reward according to $R(s,a)$. We define the set of observable histories as $\mathcal{H}=\bigcup_{t=0}^{\infty}\mathcal{H}_{t}$, where $\mathcal{H}_{t}\subseteq\mathcal{O}\times(\mathcal{A}\times\mathcal{O})^{t}$ is the set of histories of size $t$. The objective is to maximize the expected return $J(\pi)=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$, where $\pi(a\mid h)$ denotes a history-dependent policy and $\gamma\in[0,1)$ is a discount factor. $P$ specifies the initial state distribution. The history-based Q-function is defined as
$Q^{\pi}(h,a)=\mathbb{E}^{\pi}_{s_{0:\infty},a_{0:\infty}}\left[\sum_{j=0}^{\infty}\gamma^{j}R(s_{j},a_{j})\middle|H_{0}=h,A_{0}=a\right],$
and the corresponding value function as
$V^{\pi}(h)~=~\sum_{a\in\mathcal{A}}\pi(a\mid h)\ Q^{\pi}(h,a).$
In the following, we write conditional expectations by placing the conditioning in the subscript, e.g., $\mathbb{E}^{\pi}_{s_{0:\infty},\,a_{0:\infty}\mid h,a}\left[\,\cdot\,\right]$.
@@ 3.2 Informed POMDPs
The informed POMDP [19] augments the POMDP definition by a so-called information space $\mathcal{I}$ and a corresponding information function $I:\mathcal{S}\to\Delta(\mathcal{I})$, which gives the probability to obtain information $i_{t}\in\mathcal{I}$ in the true state $s_{t}\in\mathcal{S}$. Hence, the informed POMDP is defined by the 10-tuple $(\mathcal{S},\mathcal{A},\mathcal{I},\mathcal{O},T,I,\widetilde{O},R,P,\gamma)$.
In contrast to the POMDP, the observation function is defined as $\widetilde{O}:\mathcal{I}\to\Delta(\mathcal{O})$ and denotes the probability to obtain $o_{t}\in\mathcal{O}$ given $i_{t}\in\mathcal{I}$. The main assumption in an informed POMDP is that the observation $o_{t}$ is conditionally independent of the true state $s_{t}$ given the information $i_{t}$, i.e., $p(o_{t}\mid i_{t},s_{t})=\widetilde{O}(o_{t}\mid i_{t})$. Each informed POMDP induces an underlying execution POMDP defined as $(\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,P,\gamma)$, where the observation function is given by
$O(o_{t}\mid s_{t})=\sum_{i\in\mathcal{I}}\widetilde{O}(o_{t}\mid i)I(i\mid s_{t}).$
3.3 Actor-Critic Paradigm
Actor-critic methods combine a policy model (actor), parameterized by $\theta$, with a value estimator (critic), parameterized by $\vartheta$. Under partial observability, both components typically condition on the observation-action history $h_{t}\in\mathcal{H}$ and are trained via sample-based gradients. The actor selects actions using $\pi_{\theta}(a_{t}\mid h_{t})$, while the critic guides learning via value estimates.
In the symmetric setting, both actor and critic share the same input, i.e., $h_{t}$. The policy gradient is given by
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}\left[\sum_{t}\gamma^{t}Q^{\pi}(h_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid h_{t})\right].$ (1)
In practice, the critic estimates $Q^{\pi}(h_{t},a_{t})$ via the temporal-difference (TD) error
$\delta_{t}=r_{t}+\gamma\hat{V}(h_{t+1};\vartheta)-\hat{V}(h_{t};\vartheta),$
computed from value estimates $\hat{V}(\cdot;\vartheta)$.
Asymmetric actor-critic methods allow the critic to access additional information during training, unavailable to the actor and at execution. Prior approaches often rely on state-based critics $V^{\pi}(s)$ [9], which are generally ill-defined in POMDPs [10].
@@ 4 INFORMED ASYMMETRIC ACTOR-CRITIC
We introduce an asymmetric actor-critic framework that leverages arbitrary privileged information during training. Based on the informed POMDP paradigm, we define informed history-based value functions and derive an unbiased policy gradient for theoretically grounded asymmetric learning.
@@ 4.1 Informed History-based Value Functions
Given an informed POMDP $\mathcal{P}$, we first define the time-invariant informed history-based reward function $R(h,i,a)$, which incorporates additional state-conditioned information $i\sim I(i\mid s)$.
Definition 4.1 (Informed history-based reward function).
The informed history-based reward function $R(h,i,a)$ is the expected state-based reward $R(s,a)$ given the belief $p(s\mid h,i)$ about the true state $s\in\mathcal{S}$, i.e.,
$R(h,i,a)=\mathbb{E}_{s\mid h,i}\left[R(s,a)\right].$ (2)
@@
In an informed POMDP, the informed history-based reward function $R(h,i,a)$ satisfies
$\mathbb{E}_{i\mid h}\left[R(h,i,a)\right]=R(h,a),$
for all $h\in\mathcal{H}$ and $a\in\mathcal{A}$, where the expectation is taken under the belief $p(i\mid h)$.
@@
$\begin{split}R(h,a)&=\sum_{s\in\mathcal{S}}p(s\mid h)R(s,a)\\ &=\sum_{s\in\mathcal{S}}\ \Big(\sum_{i\in\mathcal{I}}p(s\mid h,i)\ p(i\mid h)\Big)R(s,a)\\ &=\sum_{i\in\mathcal{I}}\Big(\sum_{s\in\mathcal{S}}R(s,a)\ p(s\mid h,i)\Big)\ p(i\mid h)\\ &=\mathbb{E}_{i\mid h}\Big[\mathbb{E}_{s\mid h,i}\big[R(s,a)\big]\Big]=\mathbb{E}_{i\mid h}\big[R(h,i,a)\big].\end{split}$
@@
By Lemma 4.1, $R(h,i,a)$ defines an unbiased estimator of the standard history-based reward $R(h,a)$. We assume that the reward function $R(s,a)$ is uniformly bounded by a constant $r_{\max}>0$. This bound also applies to the standard and informed history-based rewards.
Next, we introduce the informed history $Q$-function, which conditions on $h$, $i$, and $a$.
Definition 4.2 (Informed history $Q$-function).
The informed history $Q$-function $Q^{\pi}(h,i,a)$ denotes the expected discounted return when starting from history $h\in\mathcal{H}$, privileged information $i\in\mathcal{I}$, and action $a\in\mathcal{A}$, and then following policy $\pi$:
$Q^{\pi}(h,i,a)=\mathbb{E}^{\pi}_{s_{0:\infty},a_{0:\infty}\mid h,i,a}\left[\sum_{j=0}^{\infty}\gamma^{j}R(s_{j},a_{j})\right].$ (4)
Lemma 4.2 (Unbiasedness of the informed $Q$-function).
@@
$\mathbb{E}_{i\mid h}\left[Q^{\pi}(h,i,a)\right]=Q^{\pi}(h,a),$
for all $h\in\mathcal{H}$ and $a\in\mathcal{A}$.
@@
$\begin{split}Q^{\pi}(h,a)&=\mathbb{E}^{\pi}_{s_{0:\infty},a_{0:\infty}\mid h,a}\Big[\sum_{j=0}^{\infty}\gamma^{j}\ R\big(s_{j},a_{j}\big)\Big]\\ &=\mathbb{E}_{i\mid h}\Big[\mathbb{E}^{\pi}_{s_{0:\infty},a_{0:\infty}\mid h,i,a}\Big[\sum_{j=0}^{\infty}\gamma^{j}\ R\big(s_{j},a_{j}\big)\Big]\Big]\\ &=\mathbb{E}_{i\mid h}\Big[Q^{\pi}(h,i,a)\Big].\end{split}$
@@
Hence, by Lemma 4.2, $Q^{\pi}(h,i,a)$ defines an unbiased estimator of the standard history-based value function $Q^{\pi}(h,a)$.
Based on the proposed $Q^{\pi}(h,i,a)$, we can define the time-invariant informed asymmetric value function that evaluates a history $h$ of past observations and actions in conjunction with state-conditioned information $i$.
Definition 4.3 (Informed history value function).
The informed value function $V^{\pi}(h,i)$ denotes the expected return starting from history $h\in\mathcal{H}$ and additional information $i\in\mathcal{I}$:
$V^{\pi}(h,i)=\mathbb{E}^{\pi}_{s_{0:\infty},a_{0:\infty}\mid h,i}\left[\sum_{j=0}^{\infty}\gamma^{j}R(s_{j},a_{j})\right].$ (5)
It satisfies the recursive form:
$V^{\pi}(h,i)=\sum_{a\in\mathcal{A}}\pi(a\mid h)\,Q^{\pi}(h,i,a),$ (6)
where the informed Q-function satisfies the Bellman equation:
$Q^{\pi}(h,i,a)=R(h,i,a)+\gamma\,\mathbb{E}_{o^{\prime},i^{\prime}\mid i,a}\left[V^{\pi}(h^{\prime},i^{\prime})\right],$ (7)
with $h^{\prime}=hao^{\prime}$.
In contrast to the history value $V^{\pi}(h)$, the informed history value $V^{\pi}(h,i)$ leverages additional state-conditioned context, potentially enabling a more comprehensive understanding of the environment’s current state and its reward structure. Similarly, unlike the state value $V^{\pi}(s)$, the history-information pair $(h,i)\in\mathcal{H}\times\mathcal{I}$ provides a richer observable basis for forecasting agent behavior.
Lemma 4.3 (Unbiasedness of the informed value function).
In an informed POMDP, the informed value function satisfies for all $h\in\mathcal{H}$:
$\mathbb{E}_{i\mid h}\left[V^{\pi}(h,i)\right]=V^{\pi}(h).$
@@
$V^{\pi}(h)=\mathbb{E}^{\pi}_{s_{0:\infty},\,a_{0:\infty}\mid h}\left[\sum_{j=0}^{\infty}\gamma^{j}R(s_{j},a_{j})\right],$
@@
$\begin{split}V^{\pi}(h)&=\mathbb{E}^{\pi}_{s_{0:\infty},a_{0:\infty}\mid h}\Big[\sum_{j}^{\infty}\gamma^{j}\ R\big(s_{j},a_{j}\big)\Big]\\ &=\mathbb{E}_{i\mid h}\Big[\mathbb{E}^{\pi}_{s_{0:\infty},a_{0:\infty}\mid h,i}\Big[\sum_{j}^{\infty}\gamma^{j}\ R\big(s_{j},a_{j}\big)\Big]\Big]\\ &=\mathbb{E}_{i\mid h}\Big[V^{\pi}(h,i)\Big].\end{split}$

### v2

@@ Abstract
Asymmetric actor-critic methods are widely used in partially observable reinforcement learning, but typically assume full state observability to condition the critic during training, which is often unrealistic in practice. We introduce the informed asymmetric actor-critic framework, allowing the critic to be conditioned on arbitrary state-dependent privileged signals without requiring access to the full state. We show that any such privileged signal yields unbiased policy gradient estimates, substantially expanding the set of admissible privileged information. This raises the problem of selecting the most adequate privileged information in order to improve learning. For this purpose, we propose two novel informativeness criteria: a dependence-based test that can be applied prior to training, and a criterion based on improvements in value prediction accuracy that can be applied post-hoc. Empirical results on partially observable benchmark tasks and synthetic environments demonstrate that carefully selected privileged signals can match or outperform full-state asymmetric baselines while relying on strictly less state information.
@@ 1 Introduction
Reinforcement learning (RL) has emerged as a powerful tool for optimizing control policies in various domains, including the control of heating, ventilation, and air conditioning systems [2], energy systems [14, 16], autonomous driving [36], and robotics [39]. However, many real-world applications involve partial observability, where agents must make decisions based on incomplete and noisy observations. This setting is formalized by partially observable Markov decision processes (POMDPs) [27], where optimal actions depend on the history of past observations and actions.
@@
To address this, RL methods for fully observable settings have been adapted by learning history-dependent policies, typically using recurrent neural networks (RNNs) that encode observation-action sequences [11, 47, 46, 23, 44, 8]. Many approaches compress these sequences into compact latent representations by introducing auxiliary learning objectives [34, 37, 32, 21, 22, 18, 20, 9, 25]. While these methods can, in principle, learn optimal policies, they assume the same level of observability during training and execution, constraining policy learning to the limited information available at deployment. Yet, in practice, this assumption is unnecessarily restrictive and possibly suboptimal. Many training environments provide additional information that is unavailable or impractical to access during execution, such as diagnostic sensors, internal simulator variables, or forecasted data. Importantly, these additional signals do not necessarily correspond to full access to the true underlying state, but they offer valuable insights that can improve training.
@@
Several approaches have explored the use of privileged information. For example, some methods train privileged expert policies conditioned on the true state and then imitate them [12]. However, these methods often lack theoretical guarantees and may lead to suboptimal policies in POMDPs [42]. To mitigate this, Warrington et al. [42] propose constraining the expert policy to yield an optimal policy under partial observability. Another line of work exploits privileged information in model-based RL by constructing world models that summarize histories or integrate additional state signals. Examples include the Informed Dreamer [30], the Wasserstein Believer [3], and the Scaffolder [24].
Asymmetric actor-critic methods offer a promising approach for leveraging privileged information by conditioning the actor on observable histories and the critic on privileged signals during training [40, 29, 13]. Performance gains therefore arise from improved value estimation rather than richer actor inputs, as the actor never observes privileged signals. Pinto et al. [35] introduce an early asymmetric actor-critic method with a state-conditioned critic that achieves strong empirical performance. However, this formulation is generally ill-defined in POMDPs unless additional assumptions hold, such as state-decodability, where the belief over latent states collapses to a Dirac distribution given the history [4]. Baisero and Amato [4] address this issue by introducing the history-state value function to ensure unbiased gradients. Theoretical work has further established convergence guarantees for policy gradient and actor-critic methods in both fully and partially observable settings, including symmetric recurrent natural actor-critic methods using RNNs [11, 1, 41] and asymmetric settings with fixed agent state and linear function approximators [31] or tabular belief-weighted formulations [10]. However, existing asymmetric actor-critic approaches often assume full-state access, leaving the case of partial privileged signals underexplored.
In this work, we generalize asymmetric actor-critic methods by introducing the informed asymmetric actor-critic framework, which allows the critic to be conditioned on arbitrary state-dependent privileged signals, instead of requiring full-state access. We show that any state-conditioned additional information can be leveraged to ensure unbiased gradients in asymmetric actor-critic methods. This generalization allows for a broader class of privileged training-time signals, significantly extending the scope of asymmetric learning. It also raises a natural question: which privileged signals should be selected?
Given that any state-conditioned signal could, in theory, yield unbiased gradients, there is no inherent reason to prefer one signal over another. Selecting and evaluating useful signals therefore becomes a central problem in the next stage of asymmetric learning. We provide a first systematic answer to this question by introducing two novel informativeness criteria that assess the utility of additional signals w.r.t. value estimation: (i) a residual-based dependency measure, and (ii) a value-prediction-based measure to evaluate the effectiveness of privileged signals. We empirically validate our framework on benchmark tasks and synthetic informed POMDPs. Our results show that leveraging informative privileged signals significantly improves policy learning, challenging the assumption that full-state access is essential for effective asymmetric RL. This work opens the door to future research on the selection, evaluation, and integration of privileged information, offering new opportunities for improving RL performance in partially observable settings.
@@ 2 Background
In this section, we formally introduce partially observable Markov decision processes (POMDPs) and the symmetric actor-critic paradigm, and then describe the informed POMDP framework that motivates our informed asymmetric actor-critic framework.
@@ 2.1 Partially Observable Markov Decision Processes
A POMDP [27] models sequential decision-making under uncertainty as tuple $(\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,P,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{O}$ denote the state, action, and observation spaces. The transition probabilities $T(s_{t+1}|s_{t},a_{t})$ describe the process dynamics. The agent emits observations via $O(o_{t}|s_{t})$ at time $t$ and selects actions $a_{t}$ based on the observable history $h_{t}$, defined as the sequence of past observations and actions, including the current observation. Specifically, the history can be written as $h_{t}=(o_{0},a_{0},\dots,o_{t})$. We define the set of observable histories as $\mathcal{H}=\bigcup_{t=0}^{\infty}\mathcal{H}_{t}$, where $\mathcal{H}_{t}\subseteq\mathcal{O}\times(\mathcal{A}\times\mathcal{O})^{t}$ is the set of histories of size $t$. The expected immediate reward of the agent is given by $R(s_{t},a_{t})$. We assume rewards are uniformly bounded by a constant $r_{\max}>0$. $P$ specifies the initial state distribution. The objective is to maximize the expected return $J(\pi)=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$, where $\pi(a_{t}|h_{t})$ denotes a history-dependent policy and $\gamma\in[0,1)$ is a discount factor. The future return from time $t$ is defined as $G_{t}~=~\sum_{k=0}^{\infty}~\gamma^{k}R(s_{t+k},a_{t+k})$, which leads to the history-based $Q$-function, the expected return given the current history and action,
$Q^{\pi}(h_{t},a_{t})=\mathbb{E}^{\pi}_{s_{t+1:\infty},a_{t+1:\infty}}\big[G_{t}\mid h_{t},a_{t}\big],$
and the corresponding history value function
$V^{\pi}(h_{t})~=~\sum_{a_{t}\in\mathcal{A}}\pi(a_{t}|h_{t})\ Q^{\pi}(h_{t},a_{t}).$
2.2 Symmetric Actor-Critic Paradigm
Given a history-dependent policy $\pi_{\theta}$ parametrized by $\theta$ (actor), the policy gradient in a POMDP is given by [38, 43]
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}\left[\sum_{t}\gamma^{t}Q^{\pi}(h_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\right].$ (1)
In practice, this expectation is estimated via Monte Carlo by sampling histories and actions in the POMDP, which is known to yield high-variance estimates. It has been shown that subtracting any function of the history, referred to as a baseline, from the $Q$-function does not bias the gradient while potentially reducing variance. Commonly, the value function $V^{\pi}(h_{t})$ is used as this baseline, yielding the advantage formulation
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}\left[\sum_{t}\gamma^{t}A^{\pi}(h_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\right],$ (2)
where $A^{\pi}(h_{t},a_{t})~=~Q^{\pi}(h_{t},a_{t})~-~V^{\pi}(h_{t})$ is called the advantage function. Unfortunately, the exact forms in Equations 1-2 involve the unknown functions $Q^{\pi}(\cdot)$ and $V^{\pi}(\cdot)$. The $Q$-function $Q^{\pi}(h_{t},a_{t})$ can in principle be estimated using a single Monte Carlo sample of $G_{t}$, but $V^{\pi}(h_{t})$ cannot because the sampled $G_{t}$ is not conditionally independent of $a_{t}$ given $h_{t}$. Practitioners typically approximate $Q^{\pi}(\cdot)$ or $V^{\pi}(\cdot)$ with parametric functions $Q^{\pi}_{\vartheta}$ or $V^{\pi}_{\vartheta}$ (critic), using temporal difference (TD) learning. Since a baseline requires $V^{\pi}(\cdot)$, many algorithms avoid estimating $Q^{\pi}$ directly and instead rely on an advantage estimator that uses only the critic, $\hat{A}_{t}~=~r_{t}~+~\gamma~V^{\pi}_{\vartheta}(h_{t+1})~-~V^{\pi}_{\vartheta}(h_{t})$, where $r_{t}$ is the observed reward at time $t$. This estimator of the advantage is unbiased if $V^{\pi}_{\vartheta}$ perfectly approximates $V^{\pi}$.
@@ 2.3 Informed POMDPs
Modeling the training-execution asymmetry, in which the critic observes additional state-dependent signals during training, but the policy at deployment relies only on observation-action histories, requires a framework that formalizes these signals without altering execution-time dynamics. The informed POMDP formalism [30] provides a principled approach by augmenting the standard POMDP with an information variable. Specifically, the informed POMDP introduces an information space $\mathcal{I}$ and an information function $I:\mathcal{S}\to\Delta(\mathcal{I})$, which gives the probability to obtain information $i_{t}\in\mathcal{I}$ in the true state $s_{t}\in\mathcal{S}$. Hence, the informed POMDP is defined by the 10-tuple $(\mathcal{S},\mathcal{A},\mathcal{I},\mathcal{O},T,I,\widetilde{O},R,P,\gamma)$. In contrast to the standard POMDP, the observation function is defined as $\widetilde{O}:\mathcal{I}\to\Delta(\mathcal{O})$ and denotes the probability to obtain $o_{t}\in\mathcal{O}$ given $i_{t}\in\mathcal{I}$. The defining assumption of the informed POMDP is that the observation $o_{t}$ is conditionally independent of the true state $s_{t}$ given the information $i_{t}$, i.e., $p(o_{t}|i_{t},s_{t})=\widetilde{O}(o_{t}|i_{t})$. The information variable can thus be interpreted as a latent mediator that captures all state-dependent information relevant for generating observations. Each informed POMDP induces an underlying execution POMDP defined as $(\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,P,\gamma)$, with observation function $O(o_{t}|s_{t})=\int_{i\in\mathcal{I}}\widetilde{O}(o_{t}|i)\,I(i|s_{t})\text{d}i$, which governs the agent’s interaction at deployment.
Remark 2.1.
[…]
Thus, the conditional-independence assumption offers a convenient modeling abstraction for introducing an information variable that is at least as informative about the underlying state as the execution-time observation, while preserving full generality. As a result, the informed POMDP framework naturally subsumes both standard POMDPs and settings with additional training-time signals, making it particularly well-suited for modeling asymmetric learning scenarios.
@@ 3 Informed Asymmetric Actor-Critic
In this section, we develop an asymmetric actor-critic framework that leverages arbitrary state-conditioned information during training. Building on the informed POMDP paradigm, we introduce informed history-based $Q$- and value functions and derive an unbiased policy gradient estimator. Our formulation strictly generalizes the history-state critic of Baisero and Amato [4]: instead of requiring access to the full state, we show that any privileged signal $i_{t}\sim I(i_{t}|s_{t})$ suffices to obtain unbiased value estimates, without imposing additional assumptions on the POMDP. This significantly broadens the class of training-time signals that can be leveraged in asymmetric actor-critic methods, and motivates the question of how such signals should be selected.
@@ 3.1 Informed History-based Value Functions
Given an informed POMDP, we consider the following informed history-based reward
$R(h_{t},i_{t},a_{t})=\mathbb{E}_{s_{t}}\big[R(s_{t},a_{t})\mid h_{t},i_{t}\big].$ (3)
This reward incorporates additional state-conditioned information available during training while remaining unbiased w.r.t. the standard history-based reward when taking expectations over $p(i_{t}|h_{t})$.
@@
In an informed POMDP, the informed history-based reward function $R(h_{t},i_{t},a_{t})$ satisfies
$\mathbb{E}_{i_{t}}\big[R(h_{t},i_{t},a_{t})|h_{t}\big]=R(h_{t},a_{t}),$
for all $h_{t}\in\mathcal{H}$ and $a_{t}\in\mathcal{A}$, where the expectation is taken under the belief $p(i_{t}|h_{t})$.
@@
$\begin{split}R(h_{t},a_{t})&=\sum_{s_{t}\in\mathcal{S}}p(s_{t}|h_{t})R(s_{t},a_{t})\\ &=\sum_{s_{t}\in\mathcal{S}}\ \Big(\sum_{i_{t}\in\mathcal{I}}p(s_{t}|h_{t},i_{t})\ p(i_{t}|h_{t})\Big)R(s_{t},a_{t})\\ &=\sum_{i_{t}\in\mathcal{I}}\Big(\sum_{s_{t}\in\mathcal{S}}R(s_{t},a_{t})\ p(s_{t}|h_{t},i_{t})\Big)\ p(i_{t}|h_{t})\\ &=\mathbb{E}_{i_{t}}\Big[\mathbb{E}_{s_{t}}\big[R(s_{t},a_{t})\big|h_{t},i_{t}\big]\Big|h_{t}\Big]\\ &=\mathbb{E}_{i_{t}}\big[R(h_{t},i_{t},a_{t})\big|h_{t}\big].\end{split}$
@@
Informed history $Q$-function. We now introduce the informed history $Q$-function, which extends the history-state critic of Baisero and Amato [4] to arbitrary state-conditioned privileged signals:
$Q^{\pi}(h_{t},i_{t},a_{t})=\mathbb{E}^{\pi}_{s_{t+1:\infty},a_{t+1:\infty}}\big[G_{t}\mid h_{t},i_{t},a_{t}\big].$
This $Q$-function is an unbiased estimator of the standard history-based $Q$-function.
Lemma 3.2 (Unbiasedness of the informed history $Q$-function).
@@
$\mathbb{E}_{i_{t}}\big[Q^{\pi}(h_{t},i_{t},a_{t})|h_{t}\big]=Q^{\pi}(h_{t},a_{t}),$
for all $h_{t}\in\mathcal{H}$ and $a_{t}\in\mathcal{A}$.
@@
$\begin{split}Q^{\pi}(h_{t},a_{t})&=\mathbb{E}^{\pi}_{s_{t+1:\infty},a_{t+1:\infty}}\Bigg[\sum_{k=0}^{\infty}\gamma^{k}\ R\big(s_{t+k},a_{t+k}\big)\Big|h_{t},a_{t}\Bigg]\\ &=\mathbb{E}_{i_{t}}\Bigg[\mathbb{E}^{\pi}_{s_{t+1:\infty},a_{t+1:\infty}}\Big[\sum_{k=0}^{\infty}\gamma^{k}\ R\big(s_{k+t},a_{k+t}\big)\big|h_{t},i_{t},a_{t}\Big]\Big|h_{t}\Bigg]\\ &=\mathbb{E}_{i_{t}}\Big[Q^{\pi}(h_{t},i_{t},a_{t})\big|h_{t}\Big].\end{split}$
@@
Informed history value function.
Based on the proposed informed history $Q$-function, we can define the informed history value function:
$V^{\pi}(h_{t},i_{t})=\mathbb{E}^{\pi}_{s_{t+1:\infty},a_{t+1:\infty}}\big[G_{t}\mid h_{t},i_{t}\big],$ (4)
which admits the time-invariant policy-based decomposition
$V^{\pi}(h,i)=\sum_{a\in\mathcal{A}}\pi(a|h)\,Q^{\pi}(h,i,a).$ (5)
The corresponding Bellman equation is
$Q^{\pi}(h,i,a)=R(h,i,a)+\gamma\,\mathbb{E}_{o^{\prime},i^{\prime}|i,a}\left[V^{\pi}(h^{\prime},i^{\prime})\right],$ (6)
where $h^{\prime}=hao^{\prime}$. Analogously to the $Q$-function case, the informed history value function is an unbiased estimator of the standard history value.
Lemma 3.3 (Unbiasedness of the informed history value function).
In an informed POMDP, the informed history value function satisfies for all $h_{t}\in\mathcal{H}$:
$\mathbb{E}_{i_{t}}\big[V^{\pi}(h_{t},i_{t})|h_{t}\big]=V^{\pi}(h_{t}).$
@@
$V^{\pi}(h_{t})=\mathbb{E}^{\pi}_{s_{t+1:\infty},\,a_{t+1:\infty}}\Bigg[\sum_{k=0}^{\infty}\gamma^{k}R(s_{k+t},a_{k+t})\big|h_{t}\Bigg],$
@@
$\begin{split}V^{\pi}(h_{t})&=\mathbb{E}^{\pi}_{s_{t+1:\infty},a_{t+1:\infty}}\Bigg[\sum_{k}^{\infty}\gamma^{k}\ R\big(s_{k+t},a_{k+t}\big)\Big|h_{t}\Bigg]\\ &=\mathbb{E}_{i_{t}}\Bigg[\mathbb{E}^{\pi}_{s_{t+1:\infty},a_{t+1:\infty}}\Big[\sum_{k}^{\infty}\gamma^{k}\ R\big(s_{k+t},a_{k+t}\big)\big|h_{t},i_{t}\Big]\Big|h_{t}\Bigg]\\ &=\mathbb{E}_{i_{t}}\Big[V^{\pi}(h_{t},i_{t})\big|h_{t}\Big].\end{split}$